OpenStudy (anonymous):

Medal and Fan A recursive function is shown below: f(1) = 4 and f(n) = f(n - 1) - 8; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? 4, -4, -12, -20, -28, ... 4, 4, 12, 20, 28, ... 4, 12, 20, 28, 36, ... 4, -12, -20, -28, -36, ...

OpenStudy (gorv):

let n=2 f(2)=f(1)-8=-4 f(3)=f(2)-8=-12 f(4)=f(3)-8=-20

OpenStudy (gorv):

first one

OpenStudy (anonymous):

thank you :)